Find the length of a rectangular lot with a perimeter of 72 m if the length is 8 m more than the width
Answer by checkley77(12844)

L=W+8
P=2L+2W
72=2(W+8)+2W
72=2W+16+2W
4W=72-16
4W=56
W=56/4
W=14 ANS. FOR THE WIDTH.
L=14+8=22 ANS. FOR THE LENGTH.
PROOF:
72=2*22+2*14
72=44+28
72=72